Consider a market whose demand and supply curves are described by the following equations: P = 40 - 2QD and P = 20 + 0.5QS. Please find the equilibrium price (P), equilibrium quantity (Q) and the price elasticity of supply (PES) at the equilibrium price.
Answer #1

Demand is P = 40 - 2Q

Supply is P = 20 + 0.5Q

At equilibrium we have

Qd = Qs

40 - 2Q = 20 + 0.5Q

20 = 2.5Q

Q = 8 units

P = $24 per unit

PES = dQ/dP * P/Q

= 2 * 24/8

= +6.00

Thus, the equilibrium price (P) is $24 per unit, equilibrium quantity (Q) is 8 units and the price elasticity of supply (PES) at the equilibrium price is +6.00.